# ShareChat: A Dataset of Chatbot Conversations in the Wild

### v2

@@ Abstract
While academic research typically treats Large Language Models (LLM) as generic text generators, they are distinct commercial products with unique interfaces and capabilities that fundamentally shape user behavior. Current datasets obscure this reality by collecting text-only data through uniform interfaces that fail to capture authentic chatbot usage. To address this limitation, we present ShareChat, a large-scale corpus of 142,808 conversations (660,293 turns) sourced directly from publicly shared URLs on ChatGPT, Perplexity, Grok, Gemini, and Claude. ShareChat distinguishes itself by preserving native platform affordances, such as citations and thinking traces, across a diverse collection covering 101 languages and the period from April 2023 to October 2025. Furthermore, ShareChat offers substantially longer context windows and greater interaction depth than prior datasets. To illustrate the dataset’s breadth, we present three case studies: a completeness analysis of intent satisfaction, a citation study of model grounding, and a temporal analysis of engagement rhythms. This work provides the community with a vital and timely resource for understanding authentic user-LLM chatbot interactions in the wild. The dataset will be publicly available.
@@ 1 Introduction
However, current research often fails to capture the complexity of real-world LLM deployment. While commercial platforms have evolved into complex ecosystems with unique features, the scientific community relies on datasets that homogenize these interactions. As seen in both synthetic xu2024magpie; li2025infinity; ding2023enhancing and real-world datasets zhao2024wildchat; zheng2023lmsys; kopf2023openassistant; zhang2025large; taori2023stanford; conover2023free, current corpora systematically strip away interface context, treating diverse products as identical, generic text boxes. For instance, WildChat zhao2024wildchat compiles about one million conversations between users and ChatGPT collected via an access gateway that offered free usage in exchange for consent to share logs. LMSYS-Chat-1M zheng2023lmsys contains one million conversations with twenty five state of the art models recorded through a single web interface (the Vicuna demo and Chatbot Arena). OpenAssistant Conversations kopf2023openassistant provides human generated assistant style dialogues and detailed preference annotations to support alignment research. These datasets are important resources with several limitations:
@@
In this study, we address these gaps by constructing a large scale, multi-platform corpus of authentic user-LLM conversations. We compile 142,808 conversations with more than 660,293 interaction turns from five widely used platforms: ChatGPT, Perplexity, Grok, Gemini, and Claude, as shown in Tabel 1. All conversations are collected from URLs that users chose to share publicly on the respective platforms, spanning more than thirty months of AI chat usage (April 2023 to October 2025) and covering 101 distinct languages. As ShareChat relies on post-hoc sharing, it therefore captures authentic usage patterns from a diverse, global user base, rather than simulated prompts or tasks designed by researchers. Each conversation record contains the complete sequence of user and assistant turns, along with platform-specific metadata, when available, timestamps, model thinking traces, or source links attached by the interface. Table 2 outlines the distinct data attributes captured across the five platforms, and the details are shown in Appendix A.2. The data collection is done under IRB approval (#28569).
@@ 2 Data Collection
For every shared conversation page, we use automated browser control with Selenium to interact with the page, retrieve the rendered HTML, and parse it into a structured JSON representation. The resulting records include the ordered sequence of conversation turns, user prompts, model responses, and platform-specific metadata. When certain elements were only visible after user interaction, we script the corresponding actions. For example, to collect the thinking content of Claude, we first locate the thinking bar element and then trigger a click on it when the content is only accessible through interaction.
@@
We prioritize user privacy by anonymizing personally identifiable information (PII). We use Microsoft’s Presidio111https://microsoft.github.io/presidio/ as the framework, Spacy honnibal2020spacy for Named Entity Recognition (NER), and custom rules to identify and remove PII across various data types, such as names, phone numbers, emails, credit cards, and URLs, in multiple languages including English, Chinese, Russian, French, Spanish, German, Portuguese, Italian, Japanese, and Korean.
@@ 3.2 Toxicity Analysis
To assess toxicity in ShareChat, we employ both Detoxify hanu2020detoxify and OpenAI Moderation API222https://platform.openai.com/docs/guides/moderation to detect toxic content in the conversations, following the approach used in WildChat zhao2024wildchat. Detoxify is a pre-trained multilingual toxicity classification model that computes toxicity scores across seven dimensions, while OpenAI Moderation provides commercial-grade content filtering. Given the language coverage limitations of Detoxify, we retain only conversational turns in languages explicitly supported by the model, with more details shown in Appendix A.4. For Detoxify, we apply the same classification threshold as WildChat, marking messages as toxic when the toxicity score exceeds 0.1, while OpenAI Moderation uses its internal flagging mechanism to identify toxic content.
@@ 3.3 Topic Analysis
We classify each user message into one of 24 predefined topic categories using multilingual LLMs Llama-3.1-8B-Instruct 333https://huggingface.co/meta-llama/Llama-3.1-8B under a few-shot prompting setup. To enable large-scale processing, we run inference with 4-bit quantization, batch size of 32. Model outputs are normalized through a post-processing pipeline to handle formatting variants and stray meta-responses, ensuring reliable mapping to our fixed taxonomy. For downstream analyses, the 24 fine-grained categories are further consolidated into seven higher-level groups to support both granular and coarse-grained comparisons across platforms. The 7 high-level categories of user requests are plotted in Figure 4.
@@ 4.2 Source Grounding Analysis
As illustrated in Figure 6, Grok exhibits a hyper-concentrated reliance on social data. Its source distribution is heavily skewed toward the X platform, which accounts for 40,624 references, nearly triple the citations for Wikipedia (12,507). This suggests a design optimized for real-time, social-media-driven context, potentially offering higher timeliness but introducing risks related to information stability and bias.
@@ A.1 Data collection
Table 3 details the target URLs used for extraction and the data collection timeframe for each platform.
| Platform | Share URL Format | Collection Period |
| ChatGPT | chatgpt.com/share/* | May 2023 to Aug 2025 |
| Perplexity | perplexity.ai/search/* | Apr 2023 to Oct 2025 |
| Grok | grok.com/share/* | Dec 2024 to Oct 2025 |
| Gemini | gemini.google.com/share/* | Apr 2024 to Sep 2025 |
| Claude | claude.ai/share/* | Not available |
@@ A.3 Turn length Distribution
A.4 Toxicity Analysis Results
Leveraging the multi-label capabilities of the Detoxify model trained on the Jigsaw Toxic Comment Classification dataset, we analyzed six distinct dimensions of harmful content: toxicity, which is a broad category for rude or disrespectful comments, severe toxicity, obscenity, threat, insult, and identity attack. After filtering conversations supported by the model, resulting in a filtered dataset of 104,107 (72.9% of the original corpus) across all five platforms. Table 4 presents the mean probability scores for each of these six dimensions by platform. The final column, “Ratio Toxic,” reports the percentage of turns where the general toxicity score exceeded the standard threshold of 0.1.
@@ A.6 Conversation Completeness Analysis
The pipeline implements a three-stage workflow adapted from the DeepEval444https://github.com/confident-ai/deepeval?tab=readme-ov-file conversation completeness metric framework. First, we extract user intentions from each conversation by prompting Qwen3-8B yang2025qwen3 with a temperature of 0.7 to identify the distinct goals or information needs expressed in the user turns, yielding a chronologically ordered list of intentions per conversation. Second, for each identified intention, we construct an evaluation prompt that presents the full conversation context alongside the specific intention and instruct Qwen3-8B to classify whether that intention was satisfied; the model produces a categorical verdict of complete (the assistant fully addressed the request), partial (the assistant provided some relevant information but the goal was incompletely met), or incomplete (the assistant failed to address the request or the conversation ended before meaningful progress). Finally, we aggregate these intention-level verdicts into a conversation-level completeness score by computing the weighted proportion of intentions, where complete verdicts contribute 1.0, partial verdicts contribute 0.5, and incomplete verdicts contribute 0.0. This three-way categorical framework allows us to capture nuanced degrees of conversational success beyond binary outcomes. We apply this pipeline uniformly across all five platforms in ShareChat, enabling direct cross-platform comparisons of how effectively different assistants satisfy user needs over the course of extended interactions. During manual coding, we observe that mis-classification often occurred when users pasted text, such as articles or messages, without explicit instructions or questions. In these instances, the LLM generate responses based solely on an inferred understanding of the content.
Figure 10 shows the conversation-level completeness scores, where ChatGPT, Claude, Grok, and Perplexity all achieve median scores of 1.0, while Gemini’s median is 0.833. Figure 11 shows the number of interactions across platforms.

### v3

@@ Abstract
While academic research typically treats Large Language Models (LLM) as generic text generators, they are distinct commercial products with unique interfaces and capabilities that fundamentally shape user behavior. Current datasets obscure this reality by collecting text-only data through uniform interfaces that fail to capture authentic chatbot usage. To address this limitation, we present ShareChat, a large-scale corpus of 142,808 conversations (660,293 turns) sourced directly from publicly shared URLs on ChatGPT, Perplexity, Grok, Gemini, and Claude. ShareChat distinguishes itself by preserving native platform affordances, such as citations and thinking traces, across a diverse collection covering 101 languages and the period from April 2023 to October 2025. Furthermore, ShareChat offers substantially longer context windows and greater interaction depth than prior datasets. To illustrate the dataset’s breadth, we present three case studies: a completeness analysis of intent satisfaction, a citation study of model grounding, and a temporal analysis of engagement rhythms. This work provides the community with a vital and timely resource for understanding authentic user-LLM chatbot interactions in the wild. The dataset is publicly available via Hugging Face 111https://github.com/raye22/ShareChat222https://huggingface.co/datasets/tucnguyen/ShareChat.
@@ 1 Introduction
However, current research often fails to capture the complexity of real-world LLM deployment. While commercial platforms have evolved into complex ecosystems with unique features, the scientific community relies on datasets that homogenize these interactions. As seen in both synthetic xu2024magpie; li2025infinity; ding2023enhancing and real-world datasets including WildChat zhao2024wildchat, LMSYS-Chat-1M zheng2023lmsys, OpenAssistant kopf2023openassistant, Alpaca taori2023stanford, ShareGPT333https://sharegpt.com/, and Dolly conover2023free, current corpora strip away interface context, treating diverse products as identical, generic text boxes. For instance, WildChat zhao2024wildchat compiles about one million conversations between users and ChatGPT collected via an access gateway that offered free usage in exchange for consent to share logs. LMSYS-Chat-1M zheng2023lmsys contains one million conversations with twenty five state of the art models recorded through a single web interface (the Vicuna demo and Chatbot Arena). OpenAssistant Conversations kopf2023openassistant provides human generated assistant style dialogues and detailed preference annotations to support alignment research. These datasets are important resources with several limitations:
@@
In this study, we address these gaps by constructing a large scale, multi-platform corpus of authentic user-LLM conversations. We compile 142,808 conversations with more than 660,293 interaction turns from five widely used platforms: ChatGPT, Perplexity, Grok, Gemini, and Claude, as shown in Tabel 1. All conversations are collected from URLs that users chose to share publicly on the respective platforms, spanning more than thirty months of AI chat usage (April 2023 to October 2025) and covering 101 distinct languages. As ShareChat relies on post-hoc sharing, it therefore captures authentic usage patterns from a diverse, global user base, rather than simulated prompts or tasks designed by researchers. Each conversation record contains the complete sequence of user and assistant turns, along with platform-specific metadata, when available, timestamps, model thinking traces, or source links attached by the interface. Table 2 outlines the distinct data attributes captured across the five platforms, and the details are shown in Appendix A.2. Data collection is done with IRB approval (#28569).
@@ 2 Data Collection
For every shared conversation page, we use an automated browser control with Selenium to interact with the page, retrieve the rendered HTML, and parse it into a structured JSON representation. The resulting records include the ordered sequence of conversation turns, user prompts, model responses, and platform-specific metadata. When certain elements were only visible after user interaction, we script the corresponding actions. For example, to collect the thinking content of Claude, we first locate the thinking bar element and then trigger a click on it when the content is only accessible through interaction.
@@
We prioritize user privacy by anonymizing personally identifiable information (PII). We use Microsoft’s Presidio444https://microsoft.github.io/presidio/ as the framework to remove PII across various data types, such as names, phone numbers, emails, credit cards, driver licenses and URLs, in multiple languages, including English, Spanish, German, French, Italian, Portuguese, Dutch, Chinese, Japanese, Russian, and Hebrew. To ensure rigorous privacy preservation, we employed an automated LLM-based evaluator to audit the PII removal process, with detailed results provided in Appendix A.9.
@@ 3.2 Toxicity Analysis
To assess toxicity in ShareChat, we employ both Detoxify hanu2020detoxify and OpenAI Moderation API555https://platform.openai.com/docs/guides/moderation to detect toxic content in the conversations, following the approach used in WildChat zhao2024wildchat. Detoxify is a pre-trained multilingual toxicity classification model that computes toxicity scores across seven dimensions, while OpenAI Moderation provides commercial-grade content filtering. Given the language coverage limitations of Detoxify, we retain only conversational turns in languages explicitly supported by the model, with more details shown in Appendix A.4. For Detoxify, we apply the same classification threshold as WildChat, marking messages as toxic when the toxicity score exceeds 0.1, while OpenAI Moderation uses its internal flagging mechanism to identify toxic content.
@@ 3.3 Topic Analysis
We classify each user message into one of 24 predefined topic categories using multilingual LLMs Llama-3.1-8B-Instruct 666https://huggingface.co/meta-llama/Llama-3.1-8B under a few-shot prompting setup. To enable large-scale processing, we run inference with 4-bit quantization, batch size of 32. Model outputs are normalized through a post-processing pipeline to handle formatting variants and stray meta-responses, ensuring reliable mapping to our fixed taxonomy. For downstream analyses, the 24 fine-grained categories are further consolidated into seven higher-level groups to support both granular and coarse-grained comparisons across platforms. The 7 high-level categories of user requests are plotted in Figure 4.
@@ 4.2 Source Grounding Analysis
As shown in Figure 6, the source of the name is extracted from the website’s top-level domain. Only top 10 of each platform are displayed in the rose graph. Grok exhibits a hyper-concentrated reliance on social data. Its source distribution is heavily skewed toward the X platform, which accounts for 40,624 references, nearly triple the citations for Wikipedia (12,507). This suggests a design optimized for real-time, social-media-driven context, potentially offering higher timeliness but introducing risks related to information stability and bias.
@@ A.1 Data collection
Table 3 details the target URLs used for extraction and the data collection timeframe for each platform. Creation timestamps were not captured for the Claude platform in this iteration; however, we plan to incorporate this metadata in future versions of the dataset.
@@ A.3 Turn length Distribution
A.4 Toxicity Analysis Results
Leveraging the multi-label capabilities of the Detoxify model trained on the Jigsaw Toxic Comment Classification dataset, we analyzed six distinct dimensions of harmful content: toxicity, which is a broad category for rude or disrespectful comments, severe toxicity, obscenity, threat, insult, and identity attack. After filtering conversations supported by the model, resulting in a filtered dataset of 104,107 (72.9% of the original corpus) across all five platforms. Table 4 presents the mean probability scores for each of these six dimensions by platform.
@@ A.6 Conversation Completeness Analysis
The pipeline implements a three-stage workflow adapted from the DeepEval777https://github.com/confident-ai/deepeval?tab=readme-ov-file conversation completeness metric framework. First, we extract user intentions from each conversation by prompting Qwen3-8B yang2025qwen3 with a temperature of 0.7 to identify the distinct goals or information needs expressed in the user turns, yielding a chronologically ordered list of intentions per conversation. Second, for each identified intention, we construct an evaluation prompt that presents the full conversation context alongside the specific intention and instruct Qwen3-8B to classify whether that intention was satisfied; the model produces a categorical verdict of complete (the assistant fully addressed the request), partial (the assistant provided some relevant information but the goal was incompletely met), or incomplete (the assistant failed to address the request or the conversation ended before meaningful progress). Finally, we aggregate these intention-level verdicts into a conversation-level completeness score by computing the weighted proportion of intentions, where complete verdicts contribute 1.0, partial verdicts contribute 0.5, and incomplete verdicts contribute 0.0. This three-way categorical framework allows us to capture nuanced degrees of conversational success beyond binary outcomes. We apply this pipeline uniformly across all five platforms in ShareChat, enabling direct cross-platform comparisons of how effectively different assistants satisfy user needs over the course of extended interactions. During manual coding, we observe that mis-classification often occurred when users pasted text, such as articles or messages, without explicit instructions or questions. In these instances, the LLM generate responses based solely on an inferred understanding of the content.
Figure 10 shows the conversation-level completeness scores, where ChatGPT, Claude, and Perplexity all achieve median scores of 1.0, while Gemini’s median is 0.833 and Grok’s median is 0.900. Figure 11 shows the number of interactions across platforms.
@@ A.8 Additional Temporal Analysis
A.9 PII Removal Validation and Data Filtering
To validate the efficacy of our de-identification pipeline, we employed GPT-OSS-120B, accessed through the Jetstream Cloud API888https://docs.jetstream-cloud.org/inference-service/overview/, as an automated auditor to detect any residual personally identifiable information in the anonymized text. The model was configured with medium reasoning effort to optimize for computational efficiency while maintaining evaluation quality. The quantitative results of this automated audit are presented in Table 7.
To further verify these automated findings, we conducted a manual review of 50 randomly sampled entries (228 turns) that the evaluator flagged as containing PII (i.e., "Positive" labels). This human verification revealed a true positive rate of only 3.7%, indicating that the vast majority of flagged instances were false positives and that the pipeline is highly effective at removing sensitive data.
Consequently, to ensure the highest privacy standards, the final released dataset excludes languages not explicitly supported by the Presidio framework. However, for conversations in these unsupported languages, we provide the original source URLs to facilitate future research while maintaining the privacy integrity of the main corpus. The final composition of the filtered dataset, utilizing strict language and PII filtering, is detailed in Table 8.
| Platform | Success Rate | Flagged | Total |
| ChatGPT | 95.20% | 51041 | 1062949 |
| Perplexity | 94.42% | 2,899 | 54,355 |
| Grok | 94.15% | 6,010 | 106,168 |
| Gemini | 95.43% | 3,302 | 72,746 |
| Claude | 97.01% | 252 | 8,504 |
| Platform | Original | Removed | % Kept |
| ChatGPT | 1,062,949 | 86,571 | 91.9% |
| Perplexity | 54,355 | 5,754 | 89.4% |
| Grok | 106,168 | 7,280 | 93.1% |
| Gemini | 72,746 | 5,078 | 93.0% |
| Claude | 8,504 | 140 | 98.4% |
| Total | 1,304,722 | 104,823 | 92.0% |